OpenStudy (anonymous):

Evaluate: 7^2logbase2x16

OpenStudy (anonymous):

\[7^{2\log_{2}16 }\] Evaluate this! :P

OpenStudy (anonymous):

What's log 16 base 2?

OpenStudy (anonymous):

is that the same as 2^x = 16 find x? (for GoBlue_)

OpenStudy (anonymous):

yes, yes it is :)

OpenStudy (anonymous):

good, cause i also have an exam tomorrow :P

OpenStudy (anonymous):

lol paulz are you from Colonel By?

OpenStudy (anonymous):

nah i wish... then i would be doing an easier exam i think are you in grade 12? if so i would have to be 7 years younger too

OpenStudy (anonymous):

is that school in canada?

OpenStudy (anonymous):

\[16=2^4\]

OpenStudy (anonymous):

\[2^{(4*2)}= x\]

OpenStudy (anonymous):

wel know this one already....::)

OpenStudy (anonymous):

\[2^8=x\]\[x=256\]

OpenStudy (anonymous):

what happen to the 7 and the log?

OpenStudy (anonymous):

CM ignored it.

OpenStudy (anonymous):

A noble strategy.

OpenStudy (anonymous):

are you suppose to ignore it?

OpenStudy (anonymous):

so the whole thing should be 7^2(log(2)(16)) or "seven to the power of all of two multiplied by log base two, sixteen close bracket" or 7^(2*4) or 7^8 or 5764801

OpenStudy (anonymous):

no you aren't supposed to ignore it. it was just a mistake.

OpenStudy (anonymous):

ok